OpenStudy (anonymous):

What is the solution to the system of equations? 5x+4y=8 x-2y=10 ( , )

OpenStudy (anonymous):

please help me

pooja195 (pooja195):

Can you use substitution here?Have you learned it?

OpenStudy (anonymous):

im not sure can u help me please

pooja195 (pooja195):

Sure so for substitution you want to isolate the x to do that subtract 2y

pooja195 (pooja195):

\[ x=2y+10\]

OpenStudy (anonymous):

so we subtract

pooja195 (pooja195):

yes to get x by itself do you get this much?

OpenStudy (anonymous):

i still dont get it if you can explian i will give you a medal

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

This is only the first step :) we want to get X alone so we can sustitute it in the other equation to find the Y value

OpenStudy (anonymous):

......

pooja195 (pooja195):

\[ 5(2y+10)+4y=8\]

OpenStudy (anonymous):

oow

OpenStudy (anonymous):

so i subtract know :/

OpenStudy (anonymous):

im new here so yeah

pooja195 (pooja195):

1)We got X by itself 2)We took the X equation and plugged it in the other problem

pooja195 (pooja195):

\(\huge\color{pink}{WELCOME~TO~OPENSTUDY!!}\)

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

welcome ok so now solve the equation i gave

OpenStudy (anonymous):

your the first person i talk too

OpenStudy (anonymous):

ok

pooja195 (pooja195):

did u solve?

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

im stuck im in the part -1y+50=80

pooja195 (pooja195):

subtract 50

pooja195 (pooja195):

from 80

OpenStudy (anonymous):

i mean 8

OpenStudy (anonymous):

its 42

OpenStudy (anonymous):

so -1y=42

OpenStudy (anonymous):

still dont get it

pooja195 (pooja195):

you should have \[14y+50=8\]

OpenStudy (anonymous):

then i did some mistake can you the the hole problem and explian it

OpenStudy (anonymous):

im not good

pooja195 (pooja195):

\[5(2y+10)+4y=8\] Distribute \[10y+50+4y=8\]

OpenStudy (anonymous):

yeah thats what i got

pooja195 (pooja195):

10y+4y=14y

OpenStudy (anonymous):

subtract 10 to 14

pooja195 (pooja195):

Thats what you did wrong your were supposed to add :)

OpenStudy (anonymous):

and you get 4

OpenStudy (anonymous):

oow

pooja195 (pooja195):

hold on let me do the problem for you

OpenStudy (anonymous):

okj thanks

pooja195 (pooja195):

\[10y+50+4y=8\]

pooja195 (pooja195):

\[10y+4y=14y\]

OpenStudy (anonymous):

soo..

pooja195 (pooja195):

\[14y+50=8\]

OpenStudy (anonymous):

whaaaa

OpenStudy (anonymous):

just keep going

pooja195 (pooja195):

subtract 50 from both sides \[14y=-42\]

pooja195 (pooja195):

\[-3\]

pooja195 (pooja195):

So y=-3 :) now to find x

OpenStudy (anonymous):

but the answer has to be some thing like this (3,0) the 0 is the one i have to find out

OpenStudy (anonymous):

what is it because i have to leave in any second by know please say

pooja195 (pooja195):

:O please post your full question

OpenStudy (anonymous):

i did

pooja195 (pooja195):

you didnt say what you need =/ or the options

OpenStudy (anonymous):

theirs no otions

pooja195 (pooja195):

im confused now @iambatman ???

OpenStudy (anonymous):

i know @rayman5766

OpenStudy (anonymous):

What's up?

OpenStudy (anonymous):

@Natalia.lovs.you

OpenStudy (anonymous):

we need your help @iambatman

OpenStudy (anonymous):

Where are you stuck

OpenStudy (anonymous):

on the question go all the way troo the top

OpenStudy (anonymous):

What method are you allowed to use

OpenStudy (anonymous):

@InfiniteMinds

OpenStudy (anonymous):

solution

OpenStudy (anonymous):

What? Quit tagging people

OpenStudy (anonymous):

Pooja was doing it right, I don't see the problem

OpenStudy (anonymous):

the answer has too bew (#,#)

OpenStudy (anonymous):

Yeah so? Put in your solved x and y solutions you got

OpenStudy (anonymous):

It's an ordered pair (x,y)

OpenStudy (anonymous):

You solve for x and y

OpenStudy (anonymous):

so can you help me

OpenStudy (anonymous):

Pooja solved for y

OpenStudy (anonymous):

Now just plug in y in any of the equations to get the following x

OpenStudy (anonymous):

can you tell me the answer because i have to leave all rrady please can u

OpenStudy (anonymous):

No, Idc if you have to leave, that's not my concern, we're not here to give you answers, just to help you with your work. If you're not interested in that, then neither am I.

pooja195 (pooja195):

\[x=2(-3)+10\] try it =/

OpenStudy (anonymous):

it pleas

OpenStudy (anonymous):

ok ill try it

OpenStudy (anonymous):

bye have too leave ill txtf oyu guys later in as other pronelm